What is y+5=1/2(x-5) in slope-intercept form?
Answer by nerdybill(7384)

"slope-intercept" form:
y = mx + b
.
So:
y+5=1/2(x-5)
y+5= (1/2)x - 5/2
y = (1/2)x - 5/2 - 5
y = (1/2)x - 5/2 - 10/2
y = (1/2)x - 15/2